2. Solve by variation of parameters:


(D ^ 2 - 2D + 1) * y = x ^ (3/2) * e ^ x

1
Expert's answer
2022-03-23T14:41:13-0400

"Homogeneous\\,\\,equation:\\\\\\left( D^2-2D+1 \\right) y=0\\\\\\lambda ^2-2\\lambda +1=0\\Rightarrow \\lambda _{1,2}=1\\Rightarrow \\\\\\Rightarrow y=C_1e^x+C_2xe^x\\\\u_1\\left( x \\right) =e^x,u_2\\left( x \\right) =xe^x\\\\y\\left( x \\right) =A\\left( x \\right) u_1\\left( x \\right) +B\\left( x \\right) u_2\\left( x \\right) \\\\\\left[ \\begin{matrix}\tu_1\\left( x \\right)&\t\tu_2\\left( x \\right)\\\\\tu_1'\\left( x \\right)&\t\tu_2'\\left( x \\right)\\\\\\end{matrix} \\right] \\left[ \\begin{array}{c}\tA'\\left( x \\right)\\\\\tB'\\left( x \\right)\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t0\\\\\tf\\left( x \\right)\\\\\\end{array} \\right] \\\\\\left[ \\begin{matrix}\te^x&\t\txe^x\\\\\te^x&\t\te^x+xe^x\\\\\\end{matrix} \\right] \\left[ \\begin{array}{c}\tA'\\left( x \\right)\\\\\tB'\\left( x \\right)\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t0\\\\\tx^{3\/2}e^x\\\\\\end{array} \\right] \\Rightarrow \\\\\\Rightarrow \\left[ \\begin{array}{c}\tA'\\left( x \\right)\\\\\tB'\\left( x \\right)\\\\\\end{array} \\right] =\\left[ \\begin{matrix}\te^x&\t\txe^x\\\\\te^x&\t\te^x+xe^x\\\\\\end{matrix} \\right] ^{-1}\\left[ \\begin{array}{c}\t0\\\\\tx^{3\/2}e^x\\\\\\end{array} \\right] =\\\\=e^{-2x}\\left[ \\begin{matrix}\te^x\\left( 1+x \\right)&\t\t-xe^x\\\\\t-e^x&\t\te^x\\\\\\end{matrix} \\right] \\left[ \\begin{array}{c}\t0\\\\\tx^{3\/2}e^x\\\\\\end{array} \\right] =\\left[ \\begin{array}{c}\t-x^{5\/2}\\\\\tx^{3\/2}\\\\\\end{array} \\right] \\\\A\\left( x \\right) =-\\frac{2}{7}x^{7\/2},B=\\frac{2}{5}x^{5\/2}\\\\y\\left( x \\right) =-\\frac{2}{7}x^{7\/2}e^x+\\frac{2}{5}x^{7\/2}e^x=-\\frac{4}{35}x^{7\/2}e^x-particular\\,\\,solution\\\\y\\left( x \\right) =-\\frac{4}{35}x^{7\/2}e^x+C_1e^x+C_2xe^x-general\\,\\,solution"